# TensoSDF: Roughness-aware Tensorial Representation for Robust Geometry and Material Reconstruction

### v1

@@ Abstract.
Reconstructing objects with realistic materials from multi-view images is problematic, since it is highly ill-posed. Although the neural reconstruction approaches have exhibited impressive reconstruction ability, they are designed for objects with specific materials (e.g., diffuse or specular materials). To this end, we propose a novel framework for robust geometry and material reconstruction, where the geometry is expressed with the implicit signed distance field (SDF) encoded by a tensorial representation, namely TensoSDF. At the core of our method is the roughness-aware incorporation of the radiance and reflectance fields, which enables a robust reconstruction of objects with arbitrary reflective materials. Furthermore, the tensorial representation enhances geometry details in the reconstructed surface and reduces the training time. Finally, we estimate the materials using an explicit mesh for efficient intersection computation and an implicit SDF for accurate representation. Consequently, our method can achieve more robust geometry reconstruction, outperform the previous works in terms of relighting quality, and reduce 50% training times and 70% inference time.
@@
${}^{\dagger}$orresponding authors.
journal: TOGjournal: TOGccs: Computing methodologies Rendering
@@ 4. Method
In this section, we present our approach, following the similar pipeline (a geometry reconstruction step and a material estimation step) as NeRO, but with several key differences, as shown in Fig. 2. We introduce a tensorial representation for SDF (Sec. 4.2) and then propose to incorporate the radiance and reflectance fields for surface reconstruction (Sec. 4.1). After reconstructing the geometry, we unify the explicit mesh and the implicit SDF field for material estimation (Sec. 4.3). Note that the incorporation of radiance and reflectance fields is a general idea, and we present it first.
@@ 4.1. Incorporating the radiance and reflectance fields
(4) $\{a,m,r\}$ $=\Theta_{\mathrm{mat}}(p,v_{f})$
(5) $L_{\mathrm{d}}$ $=\Theta_{\mathrm{direct}}(n)$
(6) $L_{\mathrm{ind}}$ $=\Theta_{\mathrm{indirect}}(p,w_{r},r)$
(7) $v$ $=\Theta_{\mathrm{occ}}(p,w_{r})$
(8) $c_{\mathrm{ref}}$ $=\rho_{\mathrm{diff}}\cdot L_{\mathrm{d}}+\rho_{\mathrm{spec}}\cdot((1-v)\cdot L_{\mathrm{d}}+v\cdot L_{\mathrm{ind}})$
where $\Theta_{\mathrm{mat}}$ is the material MLP which outputs albedo $a$, metallic $m$ and roughness $r$. $\Theta_{\mathrm{direct}}$, $\Theta_{\mathrm{indirect}}$ and $\Theta_{\mathrm{occ}}$ are the direct-light MLP, indirect-light MLP and occlusion MLP, respectively. $w_{r}$ is the reflected view direction around the normal $n$. $\rho_{\mathrm{diff}}$ and $\rho_{\mathrm{spec}}$ are the diffuse and specular terms from the microfacet BRDF (Cook and Torrance, 1982) respectively.
@@
In this way, our model can adaptively adjust the learning weights of two fields according to the surface roughness, enabling robust reconstruction of any reflective objects.
4.2. TensoSDF representation
Thanks to the roughness-aware radiance and reflectance field, we are able to reconstruct any reflective object. However, we notice that the reconstructed objects are over-smoothing, and the training time is exhausting. Hence, we introduce a new representation for the SDF – TensoSDF, combining the tensorial representation with SDF to replace the original MLP encoded ones. A well-known issue of the tensorial representation is the noisy reconstruction surface, due to the lack of global correlation. To alleviate this problem, we further introduce two smoothness priors to reduce the noise.
Tensorial representation.
(10) $V_{p}$ $=\mathrm{v}^{X}_{k}\circ\mathrm{M}^{YZ}_{k}\oplus\mathrm{v}^{Y}_{k}\circ\mathrm{M}^{XZ}_{k}\oplus\mathrm{v}^{Z}_{k}\circ\mathrm{M}^{XY}_{k}$
(11) $\{s,v_{f}\}$ $=\Theta(V_{p},p)$
Smoothness priors.
(13) $\hat{s}$ $=(1-\alpha)\cdot s+\alpha\cdot s^{\prime}$
(14) $s^{\prime}$ $=\Theta(V^{\prime}_{p},p),$
where $V^{\prime}_{p}$ is the latent vector at the top layer.
As a result, the extracted mesh is much smoother while keeping the details. Then, it is used for material estimation in the second stage.
@@ 5. Implementation details
In this section, we present our datasets, network structures, and training details. We will release our codes and datasets soon.
@@
To validate our method, we propose a new synthetic dataset consisting of six scenes with various typical material types (diffuse, glossy, and specular). We render these scenes to generate their ground-truth images, normal maps, and relighting images under five environment lights using the Cycles renderer in Blender. Moreover, we also evaluate our method on datasets from TensoIR (Jin et al., 2023) and NeRO (Liu et al., 2023), which mainly focus on diffuse or specular scenarios, respectively. For the real datasets, we use the NeILF-HDR dataset from NeILF++ (Zhang et al., 2023) and choose ACES tone mapping (Gatta et al., 2002) to obtain the low-dynamic range (LDR) datasets, since the high-dynamic range (HDR) images are unnecessary in our experiments.
@@
In the geometry reconstruction stage, we use a coarse-to-fine training strategy by setting the initial grid resolution as $128\times 128$ and increasing to the final resolution, i.e., $512\times 512$. At the first 20k steps, we only learn the reflectance field to get a rough shape and an initial roughness. Then, we learn the radiance and reflectance fields simultaneously using the roughness as the weight.
@@ 6.1. Reconstructed geometry evaluation
To evaluate the geometry reconstruction quality, we adopt the mean angular error (MAE) between the reconstructed normal and the ground-truth normal as the metric and use the Chamfer distance (CD) metric on the NeRO (Liu et al., 2023) datasets. We compare our method with NeuS (Wang et al., 2021), TensoIR (Jin et al., 2023), NeRO (Liu et al., 2023) and NeILF++ (Zhang et al., 2023). For all these methods, we use their official implementations.
In Fig. 7, we compare our method with the other four methods on our synthetic datasets. By comparison, our method can not only reconstruct the surfaces with strong reflection in all three scenes but also recover the detailed geometries (the grid in the Compressor scene and the tube in the Helmet scene). TensoIR (Jin et al., 2023) has difficulties in handling reflective surfaces, since it represents the geometry with the density field. NeRO (Liu et al., 2023) shows impressive reconstruction on simple specular objects but easily falls into local optimum for complex scenes, leading to incorrect geometries. NeuS (Wang et al., 2021) can reconstruct most non-specular surfaces except the ones with strong reflections. NeILF++ (Zhang et al., 2023) can reconstruct geometries reasonably on various materials but loses geometry details and needs a long training time (more than 12 hours for each scene). We also report the quality metrics on our synthetic datasets in Tab. 1. Our method outperforms the other four methods on the average MAE metrics across six scenes, but has slightly higher errors than NeuS on the Dragon and Motor scenes, mainly due to the noise problem.
We also conduct the comparison on the datasets from TensoIR (Jin et al., 2023) and NeRO (Liu et al., 2023), as shown in Fig. 9 and Fig. 4. Our method can handle both diffuse and specular scenes, demonstrating its robustness. Meanwhile, our method exhibits the most detailed geometry details, thanks to the TensoSDF representation.
@@ 6.2. Estimated material evaluation
Since different methods use different shading models, comparing the reconstructed materials directly is difficult. Following NeRO (Liu et al., 2023), we relight the estimated materials and compare their results with the ground truth as one effective way for validation. Specifically, we re-illuminate the scene with five new environment lights, each rendering 20 evenly distributed relighted images. We use peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) (Wang et al., 2004), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018) metrics to measure the quality of results. We compare our method against three others: TensoIR (Jin et al., 2023), NeRO (Liu et al., 2023) and NeILF++ (Zhang et al., 2023) on our synthetic dataset. We scale the relighted images with a global scalar for all the methods, as done in TensoIR (Jin et al., 2023).
@@
6.3. Results on real data
We evaluate our method on the real data from NeILF++ (Zhang et al., 2023), as shown in Fig. 8. We compare our method with NeRO (Liu et al., 2023) and NeILF++ (Zhang et al., 2023) on geometry reconstruction. Note that the input images of our method and NeRO (Liu et al., 2023) are LDR images, while NeILF++ (Zhang et al., 2023) uses the HDR images following the official implementation. By comparison, our method is able to handle surfaces with strong highlights, while NeILF++ (Zhang et al., 2023) fails (pointed by the red arrow in Qilin and Luckycat scenes). NeRO (Liu et al., 2023) can also produce reasonable results but loses many geometric details, such as the characters in the Luckycat and Brassgourd scenes. Our method produces plausible relighting results (shown in the right column of Fig. 8), demonstrating the accuracy of estimated materials.
@@ 6.3. Results on real data
Compressor 5.09 5.37 9.18 6.06 3.71
Avg. 3.39 3.54 6.18 3.85 3.00
[…]
@@
Table 3. The impact of our mesh-SDF fusion strategy for the relighting results on the Dragon and Rover scenes.
Scene mesh-SDF fusion PSNR SSIM LPIPS $\downarrow$

### v2

@@ Abstract.
Reconstructing objects with realistic materials from multi-view images is problematic, since it is highly ill-posed. Although the neural reconstruction approaches have exhibited impressive reconstruction ability, they are designed for objects with specific materials (e.g., diffuse or specular materials). To this end, we propose a novel framework for robust geometry and material reconstruction, where the geometry is expressed with the implicit signed distance field (SDF) encoded by a tensorial representation, namely TensoSDF. At the core of our method is the roughness-aware incorporation of the radiance and reflectance fields, which enables a robust reconstruction of objects with arbitrary reflective materials. Furthermore, the tensorial representation enhances geometry details in the reconstructed surface and reduces the training time. Finally, we estimate the materials using an explicit mesh for efficient intersection computation and an implicit SDF for accurate representation. Consequently, our method can achieve more robust geometry reconstruction, outperform the previous works in terms of relighting quality, and reduce 50% training times and 70% inference time. Codes and datasets are available at https://github.com/Riga2/TensoSDF.
@@
Corresponding authors.
journal: TOGjournal: TOGcopyright: acmlicensedjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 4article: 150publicationmonth: 7doi: 10.1145/3658211ccs: Computing methodologies Rendering
@@ 4. Method
In this section, we present our approach, following the similar pipeline (a geometry reconstruction step and a material estimation step) as NeRO, but with several key differences, as shown in Fig. 2. We introduce a tensorial representation for SDF (Sec. 4.1) and then propose to incorporate the radiance and reflectance fields for surface reconstruction (Sec. 4.2). After reconstructing the geometry, we unify the explicit mesh and the implicit SDF field for material estimation (Sec. 4.3).
4.1. TensoSDF representation
To represent the geometry of the object, our goal is to find an high-capacity and efficient representation. Hence, we introduce a new representation for the SDF – TensoSDF, combining the tensorial representation with SDF to replace the original MLP encoded ones. A well-known issue of the tensorial representation is the noisy reconstruction surface, due to the lack of global correlation. To alleviate this problem, we further introduce two smoothness priors to reduce the noise.
Tensorial representation.
(3) $V_{p}$ $=\mathrm{v}^{X}_{k}\circ\mathrm{M}^{YZ}_{k}\oplus\mathrm{v}^{Y}_{k}\circ\mathrm{M}^{XZ}_{k}\oplus\mathrm{v}^{Z}_{k}\circ\mathrm{M}^{XY}_{k},$
(4) $\{s,v_{f}\}$ $=\Theta(V_{p},p),$
Smoothness priors.
(6) $\hat{s}$ $=(1-\alpha)\cdot s+\alpha\cdot s^{\prime},$
(7) $s^{\prime}$ $=\Theta(V^{\prime}_{p},p),$
where $V^{\prime}_{p}$ is the latent vector at the top layer. As a result, the extracted mesh is much smoother while keeping the details.
@@ 4.2. Incorporating the radiance and reflectance fields
(9) $\{a,m,r\}$ $=\Theta_{\mathrm{mat}}(p,v_{f}),$
(10) $L_{\mathrm{d}}$ $=\Theta_{\mathrm{direct}}(n),$
(11) $L_{\mathrm{ind}}$ $=\Theta_{\mathrm{indirect}}(p,w_{r},r),$
(12) $u$ $=\Theta_{\mathrm{occ}}(p,w_{r}),$
(13) $c_{\mathrm{ref}}$ $=\rho_{\mathrm{diff}}\cdot L_{\mathrm{d}}+\rho_{\mathrm{spec}}\cdot((1-u)\cdot L_{\mathrm{d}}+u\cdot L_{\mathrm{ind}}),$
where $\Theta_{\mathrm{mat}}$ is the material MLP which outputs albedo $a$, metallic $m$ and roughness $r$. $\Theta_{\mathrm{direct}}$ and $\Theta_{\mathrm{indirect}}$ are the direct-light MLP and indirect-light MLP, respectively. $u$ is the occlusion probability from the occlusion MLP $\Theta_{\mathrm{occ}}$. $w_{r}$ is the reflected view direction around the normal $n$. $\rho_{\mathrm{diff}}$ and $\rho_{\mathrm{spec}}$ are the diffuse and specular terms from the microfacet BRDF (Cook and Torrance, 1982) respectively.
@@
The gradient of roughness $r$ is detached here for stable optimization. In this way, our model can adaptively adjust the learning weights of two fields according to the surface roughness, enabling robust reconstruction of any reflective objects.
@@ 5. Implementation details
In this section, we present our datasets, network structures, and training details.
@@
To validate our method, we propose a new synthetic dataset consisting of six scenes with various typical material types (diffuse, glossy, and specular). We render these scenes to generate their ground-truth images, normal maps, and relighting images under five environment lights using the Cycles renderer in Blender. Moreover, we also evaluate our method on datasets from TensoIR (Jin et al., 2023) and NeRO (Liu et al., 2023), which mainly focus on diffuse or specular scenarios, respectively. For the real datasets, we use the NeILF-HDR dataset from NeILF++ (Zhang et al., 2023) and Stanford-ORB dataset (Kuang et al., 2024). We adopt the ACES tone mapping (Gatta et al., 2002) to obtain the low-dynamic range (LDR) images from the NeILF-HDR dataset, since the high-dynamic range (HDR) images are unnecessary in our experiments. The Stanford-ORB dataset includes ground-truth captured meshes and relighting images, and we choose its LDR images for our experiments.
@@
During the training of the geometry reconstruction stage, we use a coarse-to-fine training strategy by setting the initial grid resolution as $128\times 128$ and increasing to the final resolution, i.e., $512\times 512$. At the first 20k steps, we only learn the reflectance field to get a rough shape and an initial roughness. Then, we compute two pixel colors from the radiance and reflectance fields. We use these two colors to compute two separate losses and combine the two losses with the roughness-aware strategy by Eqn. (2). During inference, we use the reflectance field or the recovered materials to compute the final color.
@@ 6.1. Reconstructed geometry evaluation
To evaluate the geometry reconstruction quality, we adopt the mean angular error (MAE) between the reconstructed normal and the ground-truth normal as the metric and use the Chamfer distance (CD) metric on the NeRO (Liu et al., 2023) dataset. We compare our method with NeuS (Wang et al., 2021), TensoIR (Jin et al., 2023), NeRO (Liu et al., 2023) and NeILF++ (Zhang et al., 2023). For all these methods, we use their official implementations.
In Fig. 4, we compare our method with the other four methods on our synthetic dataset. By comparison, our method can not only reconstruct the surfaces with strong reflection in all three scenes but also recover the detailed geometries (the grid in the Compressor scene and the tube in the Helmet scene). TensoIR (Jin et al., 2023) has difficulties in handling reflective surfaces, since it represents the geometry with the density field. NeRO (Liu et al., 2023) shows impressive reconstruction on simple specular objects but easily falls into local optimum for complex scenes, leading to incorrect geometries. NeuS (Wang et al., 2021) can reconstruct most non-specular surfaces except the ones with strong reflections. NeILF++ (Zhang et al., 2023) can reconstruct geometries reasonably on various materials but loses geometry details and needs a long training time (more than 12 hours for each scene). We also report the quality metrics on our synthetic dataset and the average training time in Tab. 1. Our method outperforms the other four methods on the average MAE metrics across six scenes, but has slightly higher errors than NeuS on the Dragon and Motor scenes, mainly due to the noise problem. Thanks to the tensorial representation, our method achieves shorter training times than the other three methods and equals TensoIR (Jin et al., 2023).
We also conduct the comparison on the datasets from NeRO (Liu et al., 2023) and TensoIR (Jin et al., 2023), as shown in Fig. 5 and Fig. 6. Our method can handle both diffuse and specular scenes, demonstrating its robustness. Meanwhile, our method exhibits the most detailed geometry details, thanks to the TensoSDF representation.
Figure 4. Comparison of geometry reconstruction among our method, TensoIR (Jin et al., 2023), NeRO (Liu et al., 2023), NeILF++ (Zhang et al., 2023) and NeuS (Wang et al., 2021) on our synthetic dataset.
[…]
@@ 6.2. Estimated material evaluation
Since different methods use different shading models, comparing the reconstructed materials directly is difficult. Following NeRO (Liu et al., 2023), we relight the estimated materials and compare their results with the ground truth as one effective way for validation. Specifically, we first extract the recovered meshes and materials, and re-illuminate the scene with five new environment lights in, each rendering 20 evenly distributed relighted images. We use peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) (Wang et al., 2004), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018) metrics to measure the quality of results. We compare our method against three others: TensoIR (Jin et al., 2023), NeRO (Liu et al., 2023) and NeILF++ (Zhang et al., 2023) on our synthetic dataset. We scale the relighted images with a global scalar for all the methods, as done in TensoIR (Jin et al., 2023).
@@
We visualize the estimated materials (albedo, roughness and metallic) on Helmet and Robot scenes, as shown in Fig. 8. We scale the albedo results by a global scalar, as done in TensoIR (Jin et al., 2023). By comparison, our method can achieve the closest estimated albedo maps, and reasonable metallic and roughness maps, leading to credible relighting results.
@@
Compressor 5.09 5.37 9.18 6.06 3.48
Avg. MAE 3.39 3.54 6.18 3.85 2.96
Avg. training time 6 hrs 4 hrs 8 hrs 12 hrs 4 hrs
Table 2. Relighting quality in terms of PSNR and SSIM on our synthetic dataset. Bold means the best performance and underline means the second best.
@@
6.3. Results on real data
We first evaluate our method on the real dataset from NeILF++ (Zhang et al., 2023), as shown in Fig. 9. We compare our method with NeRO (Liu et al., 2023) and NeILF++ (Zhang et al., 2023) on geometry reconstruction. Note that the input images of our method and NeRO (Liu et al., 2023) are LDR images, while NeILF++ (Zhang et al., 2023) uses the HDR images following the official implementation. By comparison, our method is able to handle surfaces with strong highlights, while NeILF++ (Zhang et al., 2023) fails (pointed by the red arrow in Qilin and Luckycat scenes). NeRO (Liu et al., 2023) can also produce reasonable results but loses many geometric details, such as the characters in the Luckycat and Brassgourd scenes. Our method produces plausible relighting results (shown in the right column of Fig. 9), demonstrating the accuracy of estimated materials.
We also evaluate our method on recent Stanford-ORB (Kuang et al., 2024) dataset. We choose five typical objects with masks. For each object, we randomly select one environment light for training and the other two lights for relighting. The quantitative measurements of the geometry reconstruction and relighting results are shown in Tab. 3. Our method outperforms NeRO (Liu et al., 2023) on all five scenes. The visual results are shown in Fig 10. NeRO (Liu et al., 2023) produces erroneous surfaces in the Teapot and Cactus scenes, and over-smoothed results in the Gnome scene. Our method can restore decent object geometries and materials, leading to plausible relighting results.
Figure 6. Comparison of geometry reconstruction among our method, NeILF++ (Zhang et al., 2023), NeRO (Liu et al., 2023) and NeuS (Wang et al., 2021) on scenes from TensoIR (Jin et al., 2023).
Figure 7. Comparison of relighting results among our method, TensoIR (Jin et al., 2023), NeILF++ (Zhang et al., 2023) and NeRO (Liu et al., 2023) on our synthetic dataset.
Figure 8. Comparison of the estimated materials among our method, NeILF++ (Zhang et al., 2023) and NeRO (Liu et al., 2023) on Helmet and Robot scenes. The PSNR metric is calculated between the estimated albedo after re-scaling and ground-truth albedo, as done in TensoIR (Jin et al., 2023).
Figure 9. Visual comparison of geometry reconstruction among our method, NeRO (Liu et al., 2023) and NeILF++ (Zhang et al., 2023) and our relighting results on the real data from NeILF++ (Zhang et al., 2023). Note that the input images of our method and NeRO (Liu et al., 2023) are LDR images, while NeILF++ (Zhang et al., 2023) uses the HDR images following the official implementation.
Figure 10. Comparison of the geometry reconstruction (Left) and relighting results (Right) between our method and NeRO (Liu et al., 2023) on Stanford-ORB (Kuang et al., 2024) dataset. The PSNR metric is averaged on two environment lights.
Figure 11. Ablation studies on the geometry reconstruction. Introducing the TensoSDF (b) can reconstruct more details than NeRO (Liu et al., 2023) (a) but with problematic surfaces. After combining the radiance field and reflectance field with a fixed balancing weight of 0.5 (c), the surface quality can be improved, but still with some flaws. Finally, using the roughness as the weight (d) achieves the best quality.
Figure 12. Ablation studies on the different balancing weights. With the radiance filed only (a), i.e., $r=1$, the reconstructed geometry shows apparent flaws on the reflective surface. With the balancing weights decrease (b-d), the reconstructed quality gradually improves. Using the roughness as the weight (e) achieves the best quality.
[…]
NeRO / Ours
PSNR
NeRO / Ours
SSIM
NeRO / Ours
[…]
NeRO / Ours
Teapot 0.395 / 0.265 26.369 / 29.971 0.978 / 0.986 0.0344 / 0.0276
Gnome 0.892 / 0.396 26.465 / 28.488 0.929 / 0.956 0.1082 / 0.0947
Cactus 4.144 / 0.199 29.081 / 31.351 0.977 / 0.984 0.0399 / 0.0362
Car 0.588 / 0.349 27.763 / 28.010 0.979 / 0.981 0.0350 / 0.0367
Grogu 4.804 / 3.968 27.081 / 30.595 0.983 / 0.990 0.0396 / 0.0362
Avg. 2.165 / 1.035 27.352 / 29.683 0.969 / 0.979 0.0514 / 0.0463
Table 4. The impact of our mesh-SDF fusion strategy for the relighting results on the Dragon and Rover scenes. Bold means the best quality.
[…]
@@ 6.3. Results on real data
6.4. Ablation study
We validate the key components of our method in Fig. 11 step by step. Starting from NeRO (Liu et al., 2023), we replace its pure MLP-based geometry representation with our TensoSDF. The TensoSDF can reconstruct more geometric details and significantly improves the quality (1.65 in MAE). Moreover, TensoSDF reduces the training time to about 50%, from 8 hours to less than 4 hours for each scene in our synthetic dataset. However, the reconstructed geometry still exhibits a noticeable dent on the surface. Then, after introducing both radiance and reflectance fields with a constant weight (0.5 in practice), the surface artifacts are alleviated. Finally, replacing the constant weight with the estimated roughness forms our complete solution and closely matches the ground truth.
To further study the effectiveness of our roughness-aware balancing strategy, we provide ablation studies on different balancing weights, as shown in Fig. 12. We first set the weight $r=1$, i.e., the radiance field only, showing apparent flaws on the reflective surfaces. With the weights decreasing and approaching the surface roughness, the reconstructed quality becomes better. When setting the balancing weight as the roughness, the quality achieves the best.
We also present the ablation of our two smoothness priors in Fig. 14. Directly using the explicit representation without any smoothness priors easily causes noise on the reconstructed geometry, especially on flat surfaces. The mipmap blending strategy of the tensor grid alleviates this problem during inference, and equipping the Gaussian smooth loss in training can produce smoother results.
Figure 14. The impact of our two smoothness priors. Reconstructing object geometry using explicit representation without any smooth priors causes noise on the smooth surface (a), while applying the mipmap blending strategy alleviates this problem (b). Finally, introducing the Gaussian smooth loss in training (c) produces the smoothest result.
6.5. Discussion and limitations
Although the high-capacity TensoSDF improves the geometric details and speeds up the training, it still introduces some limitations. One problem is the trade-off between fine details and high-frequency noise. Explicit representations are more prone to over-fitting locally, which inevitably leads to noise. A common mitigation is to use additional smooth regularization, such as our Gaussian smooth loss and mipmap blending strategy. However, smooth priors also easily cause the loss of details, as shown in Fig. 15. We leave this problem for future work.
The other problem is the trade-off between the quality and storage. Though the tensorial representation is compact compared to volume-based representations, the storage cost is higher than MLP-based methods. We analyze this trade-off on different resolutions of the tensor grid, as shown in Tab. 5. With the resolution increasing, the reconstructed geometry quality improves while the storage increases. However, even on resolution 200 $\times$ 200, our method can still produce decent results. One possible solution to reduce storage further is pruning; we leave this for future work.
Figure 15. The trade-offs between the geometric details and noise. The smooth priors can alleviate the noise problem introduced from the explicit representation (first row) while also causing the loss of details (second row).
Table 5. The trade-offs between the geometry quality and storage. Ours-200 means that the final resolution of the tensor grid is 200 $\times$ 200. With the resolution increasing, the normal MAE of the reconstructed result decreases, while the model size becomes larger. The results are from the compressor scene. Bold means the best performance and underline means the second best.
NeRO TensoIR Ours-200 Ours-300 Ours-512